Please note: You are viewing the unstyled version of this website. Either your browser does not support CSS (cascading style sheets) or it has been disabled. Skip navigation.

May 30, 2014 - ESP Update

   Print Change text to small (default) Change text to medium Change text to large

Here's the latest ESP results as of yesterday (May 29, 2014)...All 3 ESPs have just started up today, but those results will not be available until this evening.

-->ESP-1 (don) deployed near NERACOOS B was blanked showing no spots on the array for A. fundyense. There was a weak hit for the MUD2 probe.

-->ESP-2 (jake) in Casco Bay did not sample.

-->ESP-3 (dennis) deployed offshore of Pemaquid point continues to record a weak but consistent signal estimated to be in the 100-150 cells/Liter for A. fundyense range. The MUD2 probe also indicated a weakly positive signal.

We'll keep the info coming as the ESPs work stretch the 3 missions into June, all 3 ESPs are taking the weekend off,


The latest ESP results just came up to the server for today (May 30, 2014)...

All 3 ESPs continue to indicate positive signals for A. fundyense in the range of 100-200 cells/Liter.

...and ESP-2 (jake) showed a weak hit for the MUD2 probe (1 of 4 Pseudo-nitzschia sp probes on the array).

Bruce Keafer

Last updated: July 1, 2019

whoi logo

Copyright ©2007 Woods Hole Oceanographic Institution, All Rights Reserved, Privacy Policy.
Problems or questions about the site, please contact